Sodium carbonate is used to manufacture glass. It is obtained by heating sodium bicarbonate as follows: 2 NaHCO3 (s) → Na2CO3 (s) + H2O (g) + CO2 (g) ΔH = +129.3 kJ This reaction is (endothermic/exothermic). How many grams of sodium carbonate are produced if 1252 kJ of heat are used to decompose the sodium bicarbonate? g Na2CO3 (4 SF) How much heat is required to decompose 15.65 grams of NaHCO3? kJ (4 SF) Blank 1: Blank 2: Blank 3:
Answer #1

2 NaHCO3 (s) → Na2CO3 (s) + H2O (g) + CO2 (g) ΔH = +129.3 kJ

The amount of energy is required for decomposition of NaHCO3 . So in this reaction heat energy is absorbed. It is endothermic reaction . For endothermic reaction ΔH is positive .

This reaction is endothermic >>>>answer

129.3KJ energy is used for formation of 1 mole of Na2CO3

129.3KJ energy is used for formation of 106g of Na2CO3

1252Kj energy is used for formation of = 106*1252/129.3   = 1026.38g of Na2CO3 >>>>>answer

2 moles of NaHCO3 decomposes to absorbed heat energy is 129.3KJ

2*84g of NaHCO3 decomposes to absorbed heat energy is 129.3Kj

15.65g of NaHCO3 decomposes to absorbed heat energy = 129.3*15.65/(2*84)   = 12.045g of NaHCO3 >>>>answer